In a ball game a 70 kg player is running at 10 m/s when he is hit by the other player he bounces off in the opposite direction at 2 m/s. What is the players change in momentum?


1
Expert's answer
2021-05-12T11:28:19-0400

"\text{before collision:}"

"m = 70;\vec{V_1}=10;"

"\vec p_1=m*\vec{V_1} = 10*70 = 700"

"\text{after collision:}"

"m = 70;\vec{V_2}=-2;"

"\vec p_2=m*\vec{V_2} = -2*70 = -140"

"\Delta p= \vec p_2-\vec p_1= -140-700 =-840kg*\frac{m}{s}"

"\text{Answer: }\Delta p=-840kg*\frac{m}{s}"